COHERENCE RELAXATION AND ROTATIONAL REVIVALS OF THE STIMULATED AND VIRTUAL FEMTOSECOND PHOTON ECHO

\underline{IGOR PASTIRK}, VADIM V. LOZOVOY, MATTHEW COMSTOCK, and MARCOS DANTUS, Department of Chemistry, Michigan State University, East Lansing, MI, 48824.

We present experimental results using stimulated and virtual photon echo pulse sequences. The long time rang data show that stimulated photon echo cancels inhomogeneous broadening processes while the virtual echo does not. We also show an unexpected observation of ground state rotational revivals on a photon echo measurement.